# Code Like Humans: A Multi-Agent Solution for Medical Coding

### v1

@@ 5.1 End-to-end evaluation
We evaluated our implementation of the CLH framework end-to-end, comparing it directly to baseline models on MDACE. Table 1 summarizes the results. With the largest base model, the CLH approach achieves comparable performance with PLM-ICD, while handling a 70-fold larger label space.
@@
To our knowledge, no prior work has considered the full label space. Hence, when considering the realistic setting of human coders, our results can be regarded as the state-of-the-art. In the future, it may be helpful and fairer to compare medical coding systems in the full and label-constrained settings.
@@ 6.1 Are all code-able cues identified?
Figure 2 shows the retrieval recall@2525 for different ICD-10 chapters. On the x-axis, we report recall based on alphabetical terms retrieved using the evidence extractor (step 1), while the y-axis reports recall based on expert-annotated text snippets. In each case, we calculate recall for the top 25 retrieved ICD-10 terms. High recall for J00–J99 (respiratory) and N00–N99 (genitourinary) shows that both the agent and expert annotations reliably retrieve index terms for body-system diseases. When the expert annotations recall is limited, as for C00–D49 (neoplasms), S00–T88 (injury or poisoning), and V00–Y99 (external causes), the agent cannot compensate, indicating a retrieval bottleneck. In contrast, we find that for the chapters F01–F99 (mental and behavioral disorders), H00–H59 (eye and adnexa), and Z00–Z99 (factors influencing health and contact with services), the 1. evidence extractor agent underperforms significantly.
@@
Figure 2: Chapter-level comparison of retrieval recall@2525. The x-axis reports retrieving alphabetical index terms with the 1. evidence extractor, while the y-axis reports retrieval with expert-annotated evidence.
@@ 6.2 What is the impact of more candidate codes?
In this section, we study the effect of the candidate space in a controlled setting that does not rely on CHL. Concretely, at inference time we construct a per-note candidate set $C$ by combining the ground-truth codes $P=\{y_{i}:y_{i}=1\}$ with $K$ hard negatives per positive: $C=P\cup R_{K}$ with $|R_{K}|=K|P|$. We identify hard negatives by embedding the short descriptions of codes using our retriever (S-PubMedBert-MS-MARCO, see subsection D.1) and sampling the $K$ nearest neighbors that are not in $P$. This allows us to examine how the size of the candidate set affects model behavior. We then run the tabular validator (step 3) and the code reconciler (step 4) only on the candidate set $C$. The tabular validator (step 3) receives mutually exclusive choices, which simplifies selection, while the code reconciler (step 4) operates in a multi-label setting and must also decide how many codes to output.

### v2

@@ 5.1 End-to-end evaluation
We evaluated our implementation of the CLH framework end-to-end, comparing it directly to baselines on MDACE. Table 1 summarizes the results. With the largest base model, CLH achieves comparable performance with PLM-ICD, while handling a 70-fold larger label space.
@@
To our knowledge, no prior work has considered the full label space. Hence, when considering the realistic task of human coders, our results can be regarded as the state-of-the-art. In the future, it may be helpful and fairer to compare medical coding systems in the full and label-constrained settings.
@@ 6.1 Are all code-able cues identified?
Figure 2 shows the retrieval recall@$25$ for different ICD-10 chapters. X-axis shows recall based on alphabetical terms retrieved using the evidence extractor (step 1), while the y-axis reports recall based on expert-annotated text snippets. In each case, we calculate recall for the top 25 retrieved ICD-10 terms. High recall for J00–J99 (respiratory) and N00–N99 (genitourinary) shows that both the agent and expert annotations reliably retrieve index terms for body-system diseases. When the expert annotations recall is limited, as for C00–D49 (neoplasms), S00–T88 (injury or poisoning), and V00–Y99 (external causes), the agent cannot compensate, indicating a retrieval bottleneck. In contrast, we find that for the chapters F01–F99 (mental and behavioral disorders), H00–H59 (eye and adnexa), and Z00–Z99 (factors influencing health and contact with services), the evidence extractor agent underperforms significantly.
@@
Figure 2: Chapter-level comparison of retrieval recall@$25$. The x-axis reports retrieving alphabetical index terms with the 1. evidence extractor, while the y-axis reports retrieval with expert-annotated evidence.
@@ 6.2 What is the impact of more candidate codes?
In this section, we study the effect of the candidate space in a controlled setting that does not rely on CLH. At inference time we construct a per-note candidate set $C$ by combining the ground-truth codes $P=\{y_{i}:y_{i}=1\}$ with $K$ hard negatives per positive: $C=P\cup R_{K}$ with $|R_{K}|=K|P|$. We identify hard negatives by embedding the short descriptions of codes using our retriever (S-PubMedBert-MS-MARCO, see subsection D.1) and sampling the $K$ nearest neighbors that are not in $P$. This allows us to examine how the size of the candidate set affects model behavior. We then run the tabular validator (step 3) and the code reconciler (step 4) only on the candidate set $C$. The tabular validator (step 3) receives mutually exclusive choices, which simplifies selection, while the code reconciler (step 4) operates in a multi-label setting and must also decide how many codes to output.